Amy has read 220 pages in 4 days. If she maintains the same read rate, she will finish the book in a total of 9 days. How many pages is Amy's book?
Answer by lwsshak3(11628)

Amy has read 220 pages in 4 days. If she maintains the same read rate, she will finish the book in a total of 9 days. How many pages is Amy's book?
***
220+(9/4)*220=220+495=715
How many pages is Amy's book? 715